a 12cm by 5 cm rectangle is cut along its diagonal. what is the sum of the perimeters of the two triangles formed?















Answer by rothauserc(4718)

first find the hypotenuse of one of the triangles since both triangles have the same length sides.
12^2 + 5^2 = h^2
169 = h^2 and
h = sqrt(169) = 13
now the perimeter of one of the triangles is
12 + 5 + 13 = 30 cm
and the perimeter of both triangles = 2 * 30 = 60 cm